OpenStudy (sarahc):

Write the complete balanced equation for the following reaction: C7H14 + O2 CO2 + H2O

OpenStudy (sarahc):

@.Sam.

sam (.sam.):

Well first balance the carbon, C7H14 + O2 = CO2 + H2O C7H14 + O2 = 7CO2 + H2O Then there's 14 hydrogen on the LHS, So 7 for RHS C7H14 + O2 = 7CO2 + 7H2O To balance oxygen, there's 21 oxygens on the RHS, so we'll use half of those on the LHS since its O "2". \[C_7H_{14} + \frac{21}{2}O_2 = 7CO_2 + 7H_2O\] The equation is now balanced, you can multiply by 2 for both sides to eliminate fractions.

OpenStudy (sarahc):

Thank you :) mind helping with another? http://openstudy.com/study#/updates/5126dac2e4b0dbff5b3cee99